• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name ent

    Log entry time 16:13:56 on October 8,2006

    Entry number 186594

    keyword=day shift summary

    08:00 Started with Mark Stevens in Hall to work on HRS-R Dipole
    10:25 Mark out of Hall, magnet ramping up
    11:45 Almost right current, but NMR does not respond, had to reboot and struggle to get it in right setting.
    12:25 Dipole settled, NMR locked, asked for 30 muA.
    13:50 Beam problems, injector + magnet steering, waiting for experts
    14:17 Beam back while waiting for experts
    14:24 Experts in, stop run.
    15:05 Beam coming back.
    15:30 Enough statistics according to Mark's hole, move to sieve on right arm